OpenStudy (abbycross167):

Can someone please check my work? Please? I would really appreciate it... Question~ Solve the systems using the elimination method. y = x - 1 y = -x + 7

OpenStudy (abbycross167):

My Answer~ -2x+y=-1 X+y=7 = -2x+y=-1 Multiply by 1 X+y=7 multiply by 2 = 2x-1=-x+7 3x-1=7 3x=8 X=8/3 Y=13/3 Final answer~ (8/3, 13,3)

OpenStudy (abbycross167):

what my teacher said was wrong~ your first equation under Answer doesn't make sense

OpenStudy (abbycross167):

@phi could you show me what I did wrong? Please?

OpenStudy (phi):

Is this the problem y = x - 1 y = -x + 7 ?

OpenStudy (abbycross167):

Yes sir this is the problem I need help with

OpenStudy (phi):

What is the first step you have been taught to do ?

OpenStudy (abbycross167):

to Line up the x's above each other, the y's above each other, and the equals signs and the constants above each other.

OpenStudy (phi):

though you don't have to, it is best to put all equations into standard form with all the variables on the left side of the = sign for example y= x-1 I would add -x to both sides what do you get ?

OpenStudy (phi):

can you write -x on both sides ?

OpenStudy (abbycross167):

so it will look like this? y=x-x-1-x? or y=x-1-x?

OpenStudy (phi):

almost. you "do the same thing" to each side you start with y = x -1 write -x + y = x - 1 -x on the right side x -x simplifies to 0 so you get -x + y = -1

OpenStudy (phi):

notice in your answer you wrote My Answer~ -2x+y=-1 that is different from what you should get: -x + y = -1

OpenStudy (phi):

for the 2nd equation y = -x + 7 add +x to both sides

OpenStudy (abbycross167):

I am soooo sorry but I typed it wrong!! it's supposed to be ~ y = 2x - 1 y = -x + 7 NOT y = x - 1 y = -x + 7 I forgot the 2.... I am very sorry

OpenStudy (phi):

ok, then you should get -2x+y= -1 x +y = 7 if you multiply the 2nd equation by 2 (I think that is what you did next) you get -2x+ y = -1 2x+2y = 14 ------------ 0 + 3y= 13 add the 2 equations y= 13/3 and x will be 7-13/3 = 21/3 - 13/3 = 8/3 which is what you got. you did it correctly, so I am not sure what your teacher does not like

OpenStudy (abbycross167):

So does it look like I missed a step or something? @phi

OpenStudy (phi):

No, it looks ok That is why I asked what the problem was, because if you were solving the wrong problem that would explain your teacher's comment.

OpenStudy (phi):

although, looking at your answer, you solved it using "substitution" and they wanted you to solve it using "elimination"

OpenStudy (abbycross167):

ooh ok... would you mind helping me do that?

OpenStudy (phi):

You start off good, with My Answer~ -2x+y=-1 X+y=7 then you say -2x+y=-1 Multiply by 1 X+y=7 multiply by 2 (I would not bother multiplying by 1 in the 1st equation because that does not do anything) for the 2nd equation, "multiply by 2" means multiply both sides and all terms by 2 you should get 2x+2y=14 and the equations are now -2x+y= -1 2x+2y= 14

OpenStudy (phi):

the idea is when you now add the 2 equations the -2x+2x =0 (is "eliminated") so add the two equations -2x+y= -1 2x+2y= 14 can you do that ?

OpenStudy (abbycross167):

@phi I am very sorry but I fell asleep...